# Agentar-Fin-R1: Enhancing Financial Intelligence through Domain Expertise, Training Efficiency, and Advanced Reasoning

### v2

@@ 4.1 Benchmark
Table 1: Finova Dataset: Comprehensive Task Distribution
@@
Subtotal 768
Complex Reasoning Subtotal 306
@@
Total 1274
@@ 4.4 Main Results
Qwen2.5-14B-Instruct 14B 84.41 71.60 68.82 37.74 65.64 79.40 35.35 57.38 62.89
Qwen2.5-72B-Instruct 72B 87.14 76.64 74.03 48.24 71.51 82.60 43.43 63.02 68.68
DeepSeek-V3 671B 87.69 77.99 73.93 54.16 73.44 88.40 52.53 70.47 72.45
GPT-4o - 85.82 74.26 72.18 45.11 69.34 78.80 51.01 64.91 67.86
@@
Qwen3-8B 8B 84.36 76.27 73.06 54.38 72.02 93.80 59.60 76.70 73.58
Qwen3-32B 32B 88.94 80.50 78.03 59.37 76.71 95.40 63.13 79.27 77.56
Qwen-QwQ-32B 32B 89.63 82.69 81.58 61.65 78.89 93.60 61.62 77.61 78.46
@@
Qwen-Fin-R1-7B 7B 79.18 65.80 61.6 38.32 61.23 72.80 28.28 50.54 57.66
Qwen-Dianjin-R1-7B 7B 84.30 74.90 72.89 41.80 68.47 74.60 41.92 58.26 65.07
Qwen-Dianjin-R1-32B 32B 88.47 80.41 81.49 55.97 76.59 84.40 58.59 71.50 74.89
XuanYuan-FinX1-Preview 70B 81.85 69.03 70.02 50.80 67.93 71.80 42.42 57.11 64.32
Agentar-Fin-R1-8B 8B 91.52 85.09 84.24 63.55 81.10 93.40 60.10 76.75 79.65
Agentar-Fin-R1-32B 32B 92.38 87.70 86.79 69.82 84.17 93.80 68.18 80.99 83.11
@@
The comprehensive evaluation results presented in Table 2 demonstrate the superior performance of our Agentar-Fin-R1 models across both financial and general domains. Our Agentar-Fin-R1-32B model establishes a new state-of-the-art benchmark with an average score of 83.11, substantially surpassing all competing baselines. Particularly noteworthy is the consistent dominance of our models across the entire spectrum of financial evaluation benchmarks: Agentar-Fin-R1-32B achieves optimal performance on Fineva (92.38), FinEval (87.70), FinanceIQ (86.79), and Finova (69.82), while the more parameter-efficient Agentar-Fin-R1-8B variant maintains highly competitive performance despite its reduced computational footprint.
@@
Qwen2.5-14B-Instruct 14B 19.33 70.75 20.16 20.00 39.22 57.00 37.74
Qwen2.5-72B-Instruct 72B 56.67 71.11 25.97 25.00 41.67 69.00 48.24
DeepSeek-V3 671B 54.00 63.31 46.90 31.00 51.27 78.50 54.16
GPT-4o - 42.00 68.15 25.97 21.00 34.57 79.00 45.11
@@
Qwen3-8B 8B 58.00 70.83 40.31 32.00 48.12 77.00 54.38
Qwen3-32B 32B 60.67 72.20 41.09 44.00 53.78 84.50 59.37
Qwen-QwQ-32B 32B 60.00 77.88 39.53 57.00 50.50 85.00 61.65
DeepSeek-R1 671B 60.00 73.70 50.00 48.00 54.97 81.00 61.28
GPT-o1 - 54.67 80.36 51.55 40.00 52.81 83.00 60.40
@@
Qwen-Fin-R1-7B 7B 28.14 49.90 20.93 27.00 27.94 76.00 38.32
Qwen-Dianjin-R1-7B 7B 25.33 55.72 22.48 28.00 38.78 80.50 41.80
Qwen-Dianjin-R1-32B 32B 54.67 69.74 31.78 47.00 51.12 81.50 55.97
XuanYuan-FinX1-Preview 70B 49.33 69.37 29.07 48.00 36.50 72.50 50.80
Agentar-Fin-R1-8B 8B 64.00 74.19 48.06 63.00 51.54 80.50 63.55
Agentar-Fin-R1-32B 32B 66.67 86.73 53.87 69.00 55.64 87.00 69.82
@@
Agentar-Fin-R1-32B stands out with the highest overall score of 69.82, outperforming even larger-scale general-purpose models such as DeepSeek-R1 (671B parameters, 61.28) and GPT-o1 (60.40). This strong performance underscores the significant benefits of domain specialization for financial tasks, where general-purpose models fall short.
@@
In the realm of complex reasoning, which combines financial mathematics, code understanding, and intricate financial problem solving, Agentar-Fin-R1-32B leads the way with a score of 55.64. This positions our model as the best performing model in handling sophisticated financial reasoning tasks, outperforming both general-purpose models and other financial-specific models. The ability to efficiently process and solve complex reasoning tasks is vital for applications such as financial analysis, forecasting, and decision support, areas where Agentar-Fin-R1 excels due to its combination of financial expertise and advanced reasoning capabilities.
@@ 4.5.1 Ablation Study on Label System and Weighted Training Framework
Random Sampling (50% data) 86.43 79.23 76.72 58.32 92.20 58.59 75.08
Label-Only Stratified Sampling (50% data) 88.61 82.98 78.43 61.41 92.00 57.07 76.75
Ours (10% data) 88.14 81.94 77.22 60.97 93.20 58.59 76.68
Ours (30% data) 89.34 83.46 78.13 61.23 91.80 60.10 77.34
Ours (50% data) 89.97 84.24 79.91 62.81 92.60 60.10 78.10
Full Data (Vanilla SFT) 89.23 83.89 78.69 61.59 91.80 58.08 77.21
Base Model (Qwen3-8B) 84.36 76.27 73.06 54.38 93.80 59.60 73.58
@@
  • The performance progressively improves as we increase the data budget: 10% → 30% → 50% (76.68 → 77.34 → 78.10).
@@
  • Label System Impact: Label-only stratified sampling (76.75) outperforms random sampling (75.08) by 1.67, demonstrating the value of structured sampling.
  • Weighting Mechanism Impact: Our full framework (78.10) further improves upon label-only sampling by 1.35, validating the effectiveness of difficulty-aware weighting.
  • Combined Effect: The synergy between label system and weighting mechanism provides a total improvement of 3.02 over random sampling.
@@ 4.5.2 Ablation Study on Two-Stage Training Pipeline
Single-Stage (SFT Only) 90.21 84.19 82.32 62.96 94.20 58.59 78.75
Ours (Two-Stage) 91.52 85.09 84.24 63.55 93.40 59.60 79.57
Base Model (Qwen3-8B) 84.36 76.27 73.06 54.38 93.80 59.60 73.58

### v3

@@ 4.1 Benchmark
Subtotal 868
Complex Reasoning Subtotal 282
@@
Total 1350
Table 1: Finova Dataset: Comprehensive Task Distribution
@@ 4.4 Main Results
Qwen2.5-14B-Instruct 14B 84.41 71.60 68.82 37.95 65.70 79.40 35.35 57.38 62.92
Qwen2.5-72B-Instruct 72B 87.14 76.64 74.03 48.22 71.51 82.60 43.43 63.02 68.68
DeepSeek-V3 671B 87.69 77.99 73.93 54.29 73.48 88.40 52.53 70.47 72.47
GPT-4o - 85.82 74.26 72.18 45.20 69.37 78.80 51.01 64.91 67.88
@@
Qwen3-8B 8B 84.36 76.27 73.06 54.45 72.04 93.80 59.60 76.70 73.59
Qwen3-32B 32B 88.94 80.50 78.03 59.46 76.73 95.40 63.13 79.27 77.58
Qwen-QwQ-32B 32B 89.63 82.69 81.58 61.70 78.90 93.60 61.62 77.61 78.47
@@
Qwen-Fin-R1-7B 7B 79.18 65.80 61.6 38.37 61.24 72.80 28.28 50.54 57.67
Qwen-Dianjin-R1-7B 7B 84.30 74.90 72.89 41.89 68.50 74.60 41.92 58.26 65.08
Qwen-Dianjin-R1-32B 32B 88.47 80.41 81.49 56.02 76.60 84.40 58.59 71.50 74.90
XuanYuan-FinX1-Preview 70B 81.85 69.03 70.02 50.74 67.91 71.80 42.42 57.11 64.31
Agentar-Fin-R1-8B 8B 91.52 85.09 84.24 63.56 81.10 93.40 60.10 76.75 79.65
Agentar-Fin-R1-32B 32B 92.38 87.70 86.79 69.93 84.20 93.80 68.18 80.99 83.13
@@
The comprehensive evaluation results presented in Table 2 demonstrate the superior performance of our Agentar-Fin-R1 models across both financial and general domains. Our Agentar-Fin-R1-32B model establishes a new state-of-the-art benchmark with an average score of 83.13, substantially surpassing all competing baselines. Particularly noteworthy is the consistent dominance of our models across the entire spectrum of financial evaluation benchmarks: Agentar-Fin-R1-32B achieves optimal performance on Fineva (92.38), FinEval (87.70), FinanceIQ (86.79), and Finova (69.93), while the more parameter-efficient Agentar-Fin-R1-8B variant maintains highly competitive performance despite its reduced computational footprint.
@@
Qwen2.5-14B-Instruct 14B 19.33 70.75 20.16 20.00 40.74 57.00 37.95
Qwen2.5-72B-Instruct 72B 56.67 71.11 25.97 25.00 41.57 69.00 48.22
DeepSeek-V3 671B 54.00 63.31 46.90 31.00 52.05 78.50 54.29
GPT-4o - 42.00 68.15 25.97 21.00 35.10 79.00 45.20
@@
Qwen3-8B 8B 58.00 70.83 40.31 32.00 48.57 77.00 54.45
Qwen3-32B 32B 60.67 72.20 41.09 44.00 54.29 84.50 59.46
Qwen-QwQ-32B 32B 60.00 77.88 39.53 57.00 50.78 85.00 61.70
DeepSeek-R1 671B 60.00 73.70 50.00 48.00 54.96 81.00 61.28
GPT-o1 - 54.67 80.36 51.55 40.00 53.19 83.00 60.46
@@
Qwen-Fin-R1-7B 7B 28.14 49.90 20.93 27.00 28.22 76.00 38.37
Qwen-Dianjin-R1-7B 7B 25.33 55.72 22.48 28.00 39.28 80.50 41.89
Qwen-Dianjin-R1-32B 32B 54.67 69.74 31.78 47.00 51.44 81.50 56.02
XuanYuan-FinX1-Preview 70B 49.33 69.37 29.07 48.00 36.14 72.50 50.74
Agentar-Fin-R1-8B 8B 64.00 74.19 48.06 63.00 51.63 80.50 63.56
Agentar-Fin-R1-32B 32B 66.67 86.73 53.87 69.00 56.33 87.00 69.93
@@
Agentar-Fin-R1-32B stands out with the highest overall score of 69.93, outperforming even larger-scale general-purpose models such as DeepSeek-R1 (671B parameters, 61.28) and GPT-o1 (60.46). This strong performance underscores the significant benefits of domain specialization for financial tasks, where general-purpose models fall short.
@@
In the realm of complex reasoning, which combines financial mathematics, code understanding, and intricate financial problem solving, Agentar-Fin-R1-32B leads the way with a score of 56.33. This positions our model as the best performing model in handling sophisticated financial reasoning tasks, outperforming both general-purpose models and other financial-specific models. The ability to efficiently process and solve complex reasoning tasks is vital for applications such as financial analysis, forecasting, and decision support, areas where Agentar-Fin-R1 excels due to its combination of financial expertise and advanced reasoning capabilities.
@@ 4.5.1 Ablation Study on Label System and Weighted Training Framework
Random Sampling (50% data) 86.43 79.23 76.72 58.73 92.20 58.59 75.32
Label-Only Stratified Sampling (50% data) 88.61 82.98 78.43 61.32 92.00 57.07 76.74
Ours (10% data) 88.14 81.94 77.22 61.01 93.20 58.59 76.68
Ours (30% data) 89.34 83.46 78.13 61.28 91.80 60.10 77.35
Ours (50% data) 89.97 84.24 79.91 62.92 92.60 60.10 78.12
Full Data (Vanilla SFT) 89.23 83.89 78.69 61.63 91.80 58.08 77.22
Base Model (Qwen3-8B) 84.36 76.27 73.06 54.45 93.80 59.60 73.59
@@
  • The performance progressively improves as we increase the data budget: 10% → 30% → 50% (76.68 → 77.35 → 78.12).
@@
  • Label System Impact: Label-only stratified sampling (76.74) outperforms random sampling (75.32) by 1.42, demonstrating the value of structured sampling.
  • Weighting Mechanism Impact: Our full framework (78.12) further improves upon label-only sampling by 1.38, validating the effectiveness of difficulty-aware weighting.
  • Combined Effect: The synergy between label system and weighting mechanism provides a total improvement of 2.80 over random sampling.
@@ 4.5.2 Ablation Study on Two-Stage Training Pipeline
Single-Stage (SFT Only) 90.21 84.19 82.32 62.87 94.20 58.59 78.73
Ours (Two-Stage) 91.52 85.09 84.24 63.56 93.40 59.60 79.57
Base Model (Qwen3-8B) 84.36 76.27 73.06 54.45 93.80 59.60 73.59